If Penny bought a stock for $80 dollars and could sell it 15 years later for 4 times what she originally paid, what is Penny’s return on owning this stock? (Enter your answer as a whole percentage (e.g. .35 should be entered as 35).)

What is Penny's return on stock ownership % ?
Answer #1

Sol :

Returnon stock =

(Ending stock price - Initial stock price + Dividends) / Initial stock price

= (320 - 80 + 0) / 80

= 240 / 80

= 3

Return on stock as percentage = 3 * 100

= 300%
Answer #2

To calculate Penny's return on owning the stock, we need to determine the percentage increase in the value of the stock over the 15-year period.

Penny bought the stock for $80 and sold it 15 years later for 4 times what she originally paid. This means the selling price is $80 * 4 = $320.

The increase in value is the difference between the selling price and the purchase price: $320 - $80 = $240.

To calculate the percentage increase, we divide the increase by the original purchase price and multiply by 100:

Percentage increase = (increase / original price) * 100 = ($240 / $80) * 100 = 300%

Therefore, Penny's return on owning this stock is 300%.


answered by: mervetokaz